An electron is projected at an angle of 28.8° above the horizontal at a speed of 8.36E+5 m/s in a region where the electric field is E = 395j N/C. Neglecting the effects of gravity, 1,calculate the time it takes the electron to return to its initial height.

2.Calculate the maximum height it reaches. Measure the height from where the electron enters the region of electric field.
Answer #1

acceleration: a= 395*e/m_e   = 395*1.6*10^-19/9.1 × 10-31 = 6.945*10^13 m/s^2

the time t is calculated as follows:

t=2vsin@/a = 1.16*10^-8 sec
b)

maximum height: h=(vsin@)^2/(2a) =1.17*10^-3 m